The Gilster Company, a machine tooling firm, has several plants. One plant, located in St. Falls, Minnesota, uses a job order costing system for its batch production processes. The St. Falls plant has two departments through which most jobs pass. Plantwide overhead, which includes the plant manager’s salary, accounting personnel, cafeteria, and human resources, is budgeted at $250,000. During the past year, actual plantwide overhead was $240,000. Each department’s overhead consists primarily of depreciation and other machine-related expenses. Selected budgeted and actual data from the St. Falls plant for the past year are as follows.

Department A Department B
Budgeted department overhead
(excludes plantwide overhead) $ 150,000 $ 600,000
Actual department overhead 160,000 620,000
Expected total activity:
Direct labor hours 35,000 15,000
Machine-hours 10,000 40,000
Actual activity:
Direct labor hours 51,000 9,000
Machine-hours 10,500 42,000

For the coming year, the accountants at St. Falls are in the process of helping the sales force create bids for several jobs. Projected data pertaining only to job no. 110 are as follows.

Direct materials $ 25,000
Direct labor cost:
Department A (2,200 hr) 45,000
Department B (800 hr) 10,000
Machine-hours projected:
Department A 200
Department B 1,200
Units produced 10,000

Assume the St. Falls plant uses three separate overhead rates to assign overhead costs to jobs.

b-1. Find the plant wide overhead rate by using expected machine hours.

b-2. Find the department overhead rate using expected machine hours for Department A and Department B.

b-3. Calculate the projected manufacturing costs for job 110 using the three separate rates computed in b-1 and b-2
Answer #1

B1

Plantwide overhead = $250000

Plant wide machine hours = 10000+40000 = 50000

Plant wide overhead rate = $250000/50000 = $5 per machine hour

B2.

Department overhead rate

Department A = $150000/10000 = $15 per machine hour

Department B = $600000/40000 = $15 per machines hour

B3 calculation of manufacturing cost of job 110

Particular amount
Material $25000
Labor
Department A $45000
Department B $10000
Overhead
Plant wide (200+1200)*$5 $7000
Department A (200*$15) $3000
Department B(1200*$15) $18000
Total cost $108000

Feel free to ask any queries..

Also plz upvote it means a lot .. thank you